# Four-dimensional phase space tomography from one-dimensional measurements of a hadron beam

### v2

@@ III.1 Accelerator and diagnostics
Our experiments were conducted at the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory (ORNL), shown in Fig. 1. The SNS bombards a liquid mercury target with intense proton pulses at 60 Hz repetition rate [9]. Each pulse is the superposition of 1000 “minipulses” injected from a 400-meter linac into a 250-meter storage ring at 1.3 GeV kinetic energy. The final pulse contains approximately $1.4\times 10^{14}$ protons, corresponding to 1.7 MW beam power. The accumulated beam is extracted and transported from the ring to the target in a 150-meter Ring-Target Beam Transport (RTBT).
@@ III.3 Beam accumulation
We measured the beam during an initial test of the eigenpainting method. We will not describe the details of this experiment here; the important point is that we intended to generate cross-plane correlations in the beam. We measured the beam for the three sets of optics described earlier, generating 36 profiles. We approximated the accelerator transport as a linear transformation of the phase space coordinates and computed the transport matrices using the OpenXAL [14] online model of the SNS. (Appendix A reports a measurement of a production beam; however, we were only able to collect two sets of optics (24 profiles) in this experiment.)
@@ V.1 Modeling and measurement errors
with $\bm{\mathbf{M}}_{xy}=\bm{\mathbf{M}}_{yx}=\bm{\mathbf{0}}$. The following are possible errors in this model: (i) The true quadruple coefficients may not be equal to the readback values. We assume these values are accurate because measured twiss parameters typically agree with design values [10]; future experiments could validate the linear model using orbit-difference measurements. (ii) Tilted quadrupoles would generate off-block-diagonal terms in the transfer matrix. We expect skew quadrupole components to be negligible, as they can be measured using BPM data [16] and would result in a tilted image on the target during neutron production. (iii) An incorrect beam energy would cause a systematic error in the quadrupole focusing strengths. We ignore this source of error because the beam energy is known to within 0.1% from time-of-flight measurements in the ring [17]. (iv) The quadrupole coefficients depend on the particle energy—each particle sees a slightly different force—but our model ignores energy spread. Strong transverse-longitudinal correlations in the bunch could amplify the error. The energy spread in the SNS is only a few MeV on top of a 1 GeV synchronous particle energy, so we expect that our model is sufficient. Including energy-dependent focusing would require a six-dimensional reconstruction. (v) The model does not include space charge. The tune shift over one turn in the ring is approximately 3%, too small to make a difference in the RTBT wire scanner region [10]. (vi) The wire scanner measurements assume the distribution has no pulse-to-pulse variation. Back-to-back measurements show almost no change in the profiles. In the future, an electron scanner [18] could validate this assumption by measuring the full profile on each pulse.
@@ Appendix A Measurement of a production beam
This appendix reports the measurement of a production beam in the SNS. By “production beam”, we mean that the accelerator was left in its default state, tuned for losses during neutron production at 1.7 MW beam power at 1.3 GeV beam energy. Fig. 9 shows the reconstructed distribution.

### v3

@@ III.1 Accelerator and diagnostics
Our experiments were conducted at the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory (ORNL), shown in Fig. 1. The SNS bombards a liquid mercury target with intense proton pulses at 60 Hz repetition rate [9]. Each pulse is the superposition of 1000 “minipulses” injected from a 400-meter linac into a 250-meter storage ring at 1 GeV kinetic energy. The final pulse contains approximately $1.4\times 10^{14}$ protons, corresponding to 1.4 MW beam power  [^1] These numbers correspond to the date of our experiments. The current beam power at the SNS is 1.7 MW and will soon increase to 2.4 MW.. The accumulated beam is extracted and transported from the ring to the target in a 150-meter Ring-Target Beam Transport (RTBT).
@@ III.3 Beam accumulation
We measured the beam during an initial test of the eigenpainting method. This method required a lower beam energy of 0.8 GeV to boost the effective strength of the injection kicker magnets. We will not describe the details of this experiment here; the important point is that we intended to generate cross-plane correlations in the beam. We measured the beam for the three sets of optics described earlier, generating 36 profiles. We approximated the accelerator transport as a linear transformation of the phase space coordinates and computed the transport matrices using the OpenXAL [15] online model of the SNS. (Appendix A reports a measurement of a production beam; however, we were only able to collect two sets of optics (24 profiles) in this experiment.)
@@ V.1 Modeling and measurement errors
with $\bm{\mathbf{M}}_{xy}=\bm{\mathbf{M}}_{yx}=\bm{\mathbf{0}}$. The following are possible errors in this model: (i) The true quadruple coefficients may not be equal to the readback values. We assume these values are accurate because measured twiss parameters typically agree with design values [11]; future experiments could validate the linear model using orbit-difference measurements. (ii) Tilted quadrupoles would generate off-block-diagonal terms in the transfer matrix. We expect skew quadrupole components to be negligible, as they can be measured using BPM data [17] and would result in a tilted image on the target during neutron production. (iii) An incorrect beam energy would cause a systematic error in the quadrupole focusing strengths. We ignore this source of error because the beam energy is known to within 0.1% from time-of-flight measurements in the ring [18]. (iv) The quadrupole coefficients depend on the particle energy—each particle sees a slightly different force—but our model ignores energy spread. Strong transverse-longitudinal correlations in the bunch could amplify the error. The nominal energy spread in the SNS is only a few MeV on top of a 1 GeV synchronous particle energy, so we expect that our model is sufficient  [^3] The beam energy in this experiment was 0.8 GeV, leading to a slightly larger fractional energy spread.. Including energy-dependent focusing would require a six-dimensional reconstruction. (v) The model does not include space charge. The tune shift over one turn in the ring is approximately 3%, too small to make a difference in the RTBT wire scanner region [11]. (vi) The wire scanner measurements assume the distribution has no pulse-to-pulse variation. Back-to-back measurements show almost no change in the profiles. In the future, an electron scanner [20] could validate this assumption by measuring the full profile on each pulse.
@@ Appendix A Measurement of a production beam
This appendix reports the measurement of a production beam in the SNS. By “production beam”, we mean that the accelerator was left in its default state, tuned for losses during neutron production at 1.4 MW beam power. Fig. 9 shows the reconstructed distribution.